OpenStudy (anonymous):

A car starts from rest and travels 60 m/s in 4 seconds. What is the acceleration? Show formula, work, and units for each variable.

OpenStudy (anonymous):

Acceleration = m/s^2 = m/s/s (60m/s)/(4s)=15m/s^2